# Supercell Wannier functions and faithful low-energy model for Bernal bilayer graphene

### v1

@@ Discussion.
In this Letter, we introduced SWFs as an efficient representation of the low-energy physics of BBG and related rhombohedral graphene stacks at low electronic densities. The microscopic SWFs bridge between the atomic-scale lattice of carbon atoms and the characteristic mesoscopic length scale at which flavor-polarized ‘wave packets’ shape the emergent physics of BBG at low electronic densities. We demonstrate that the spectral weight and Berry curvature of the low-energy Bloch states are faithfully represented by two valley-polarized $p_{\pm}$-orbitals located on the triangular sites of the superlattice. The SFW construction remains valid as long as the superlattice length scale $L_{s}=n_{s}|\bm{a}_{i}|\lesssim 1/|\bm{q}_{\text{FS}}|$ does not exceed the length scale associated to the width of the Fermi pockets $|\bm{q}_{\mathrm{FS}}|$ at a given electronic density. Otherwise, additional bands are back-folded to the mini-BZ and need to be wannierized in order to sufficiently model the low-energy physics. The ‘three-patch’ continuum model [78] proposed in the context of studying the translational symmetry-breaking anomalous Hall crystal phase [1, 75, 79, 78] can in fact be understood as special case of our microscopic SWF construction being valid in the intermediate regime $L_{s}\sim 1/|\bm{q}_{\text{FS}}|$, where exactly three bands contribute to the Fermi surface. Since valley conservation is included in the SWF model as an emergent symmetry, valley orders are encoded as on-site order parameters, as opposed to the microscopic picture, where the valley operator acts on bonds  [^2] Using valley as a quantum number in the supercell model amounts to convoluting the $\bm{q}=0$ valley operator with a $\bm{q}=\bm{G}_{s}$ charge density wave, see Supplementary Material.. This natural representation of valley as a quantum number in conjunction with the substantially reduced Hilbert space of the minimal SWF model enables the study of competing correlated phenomena in the phase diagram of rhombohedral graphene stacks with unbiased weak-coupling methods like FRG [50, 51, 81] from first-principles. To this end, it may be advisable to improve the downfolding scheme of the bare, dual-gated Coulomb interactions by using constrained random-phase approximation [82] or FRG [83] schemes. Due to its strength in representing first-principle interactions, the SWF model should allow to answer pressing questions regarding the formation of superconductivity from repulsive electron-electron or attractive phonon-induced interactions in various multilayer graphene stacks. Moreover, effects like proximity-induced Ising-SOC, microscopic defects and strain can readily be included on the level of the atomistic supercell, which allows for a material-realistic description relevant for experimental setups. In addition, the SWF procedure can be extended to other graphene multilayer stacks [57, 26, 58, 84, 85], or modulation-doped AlAs quantum wells [59, 60] featuring valley-polarized and nematic behavior at low electronic densities.
@@ Acknowledgements.
We thank Matthew Bunney, Francesco Grandi, and Jonas B. Profe for fruitful discussions. This work was supported by the Excellence Initiative of the German federal and state governments, the Ministry of Innovation of North Rhine-Westphalia and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). AF and DMK acknowledge funding by the DFG under RTG 1995, within the Priority Program SPP 2244 “2DMP” — 443273985. LK and TOW greatfully acknowledge support from the DFG through FOR 5249 (QUAST, Project No. 449872909) and SPP 2244 (Project No. 422707584). TOW is supported by the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the DFG (EXC 2056, Project ID 390715994). DMK acknowledges support by the Max Planck-New York City Center for Nonequilibrium Quantum Phenomena. We acknowledge computational resources provided by RWTH Aachen University under project numbers rwth1420.

### v2

@@ Discussion.
In this Letter, we introduced SWFs as an efficient representation of the low-energy physics of BBG and related rhombohedral graphene stacks. The Wannier construction is based on the atomic-scale lattice of carbon atoms, but describes the emergent low-energy physics in terms of localized, flavor-polarized supercell orbitals that have the emergent symmetries of continuum models imprinted, i.e., valley-$U(1)$, spin-$SU(2)$ and time-reversal symmetry $\mathcal{T}$. To this end, we demonstrate that the spectral weight and Berry curvature of the low-energy Bloch states are faithfully represented by two valley-polarized $p_{\pm}$-orbitals located on the triangular sites of the superlattice. The local nature of our model implies that long-ranged Coulomb interactions can be cast into fairly local interactions, whose strength can be controlled by metallic gates. The values of the effective Coulomb interactions in the SWF basis are estimated from first principles, including density-density and exchange-type interactions like inter-valley Hund’s couplings terms. The SFW construction remains valid as long as the superlattice length scale $L_{s}=n_{s}|\bm{a}_{i}|\lesssim 1/|\bm{q}_{\text{FS}}|$ does not exceed the length scale associated to the width of the Fermi pockets $|\bm{q}_{\mathrm{FS}}|$ at a given electronic density. Otherwise, additional bands are back-folded to the mini-BZ and need to be wannierized in order to sufficiently model the low-energy physics. The ‘three-patch’ continuum model [84] proposed in the context of studying the translational symmetry-breaking anomalous Hall crystal phase [1, 82, 85, 84] can in fact be understood as special case of our microscopic SWF construction being valid in the intermediate regime $L_{s}\sim 1/|\bm{q}_{\text{FS}}|$, where exactly three bands contribute to the Fermi surface. Since valley conservation is included in the SWF model as an emergent symmetry, valley orders are encoded as on-site order parameters, as opposed to the microscopic picture, where the valley operator acts on bonds  [^2] Using valley as a quantum number in the supercell model amounts to convoluting the $\bm{q}=0$ valley operator with a $\bm{q}=\bm{G}_{s}$ charge density wave, see SM [75].. This natural representation of valley as a quantum number in conjunction with the substantially reduced Hilbert space of the minimal SWF model enables the study of competing correlated phenomena in the phase diagram of rhombohedral graphene stacks with unbiased weak-coupling methods like FRG [50, 51, 87] from first-principles. To this end, it may be advisable to improve the downfolding scheme of the bare, dual-gated Coulomb interactions by using constrained random-phase approximation [88] or FRG [89] schemes. Due to its strength in representing first-principle interactions, the SWF model should allow to answer pressing questions regarding the formation of superconductivity from repulsive electron-electron or attractive phonon-induced interactions in various multilayer graphene stacks. Moreover, effects like proximity-induced Ising-SOC, microscopic defects and strain can readily be included on the level of the atomistic supercell, which allows for a material-realistic description relevant for experimental setups. In addition, the SWF procedure can be extended to other graphene multilayer stacks [57, 26, 58, 90, 91], or modulation-doped AlAs quantum wells [59, 60] featuring valley-polarized and nematic behavior at low electronic densities.
@@ Acknowledgements.
We thank Matthew T. Bunney, Francesco Grandi, and Jonas B. Profe for fruitful discussions. This work was supported by the Excellence Initiative of the German federal and state governments, the Ministry of Innovation of North Rhine-Westphalia and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). AF and DMK acknowledge funding by the DFG under RTG 1995, within the Priority Program SPP 2244 “2DMP” – 443273985. LK and TOW greatfully acknowledge support from the DFG through FOR 5249 (QUAST, Project No. 449872909) and SPP 2244 (Project No. 422707584). TOW is supported by the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the DFG (EXC 2056, Project ID 390715994). DMK acknowledges support by the Max Planck-New York City Center for Nonequilibrium Quantum Phenomena. We acknowledge computational resources provided by RWTH Aachen University under project numbers rwth1420.